OpenStudy (anonymous):

whats the domain and range of y=x^2-9 in interval notation

OpenStudy (anonymous):

@Hero if you could help it would be appreciated

hero (hero):

Are there any restrictions for the domain? In other words, is there a value of x that we can put into the equation that will result in a false or undefined value for y?

OpenStudy (anonymous):

yes it would be three and then would we put it in the round bracket system or is there another way for the domain

hero (hero):

If we put 3 in place of x, what would be the false value of y?

OpenStudy (anonymous):

0

hero (hero):

So when x = 3, y = 0 which means y is false? So if y = 0, that means y is false?

OpenStudy (anonymous):

yes

hero (hero):

No, y can equal zero. Just because y = 0 does not mean the value is false. y can equal zero.

OpenStudy (anonymous):

so would it be all real numbers (-infinity,infinity)

hero (hero):

The domain is all real numbers, correct. What about the range? Are there any restrictions on the domain?

OpenStudy (anonymous):

no

hero (hero):

Are you sure about that? What happens when y = -1? Can we solve for x?

OpenStudy (jdoe0001):

\(\bf y=x^2-9 \implies y=(x-{\color{brown}{ 0}})^2{\color{brown}{ -9}}\)

OpenStudy (anonymous):

no

hero (hero):

Actually, we are able to solve for x when y = -1.

OpenStudy (anonymous):

if x is has a decimal value also i guess

hero (hero):

If y = -1 then: -1 = x^2 - 9 9 - 1 = x^2 8 = x^2 x = + or - sqrt(8)

hero (hero):

But what about if y = -10?

OpenStudy (anonymous):

x would be undefined

OpenStudy (anonymous):

or would it be 1i

OpenStudy (anonymous):

i mean i could we use imaginary numbers

hero (hero):

When it comes to domain and range, by default, it is implied that we are speaking in the realm of real numbers unless otherwise specified.

hero (hero):

What we know at this point is that y = -10 is undefined, but y = -9 is defined. Knowing this, what is the range of y?

OpenStudy (anonymous):

(-10,infinity)

hero (hero):

Range -> [-9, infinity)

OpenStudy (anonymous):

ya i see what i did wrong thank you again for your help i appreciate the time you take to assist me and hope you have a good day @Hero